# Hyperelliptic Shimura curves and $L$-functions of central vanishing order at least 3

### v1

@@ 5. Triple product $L$-functions of central vanishing order at least 3
We keep the notations from the last section, except that from now on, the notation $\pi$ or $\pi_{i}$ will denote an automorphic representation of $B^{\times}$, instead of its finite part. We always assume that it is holomorphic of weight $2$ in the sense that its Jacquet–Langlands correspondence to ${\mathrm{GL}}_{2,F}$ is a cuspidal automorphic representation holomorphic of weight $2$.
@@
Here the product is over all places of $F$. Let $\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})$ be the local root number of $\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v}$. A result of Prasad [18, 19] states the following.
@@
then it has dimension 1 and $\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=\epsilon(B_{v})$.
@@
for all finite places $v$ of $F$, then $\epsilon(\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
@@
At every infinite place $v$, $\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-1$ (see [18, Section 9]). At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. So by 5.1, we have
$\epsilon(\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=\prod_{v}\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-\prod_{v}\epsilon(B_{v})=-1.$
Let $S_{B}$ be the finite set of finite places $v$ of $F$ such that $B_{v}$ is a division quaternion algebra. Below we only consider a maximal open compact subgroup $U=\prod\limits_{v\text{ finite}}U_{v}\subset PB^{\times}({\mathbb{A}}_{\mathrm{f}})$. In other words, its $v$-component is
@@
where ${\mathfrak{m}}_{F_{v}}$ is the maximal ideal of ${\mathcal{O}}_{F_{v}}$ and
(5.3) $g_{v}\in\begin{bmatrix}0&1\\ {\mathcal{O}}_{F_{v}}\backslash{\mathfrak{m}}_{F_{v}}&0.\end{bmatrix}$
If $\pi^{U}\neq 0$, then $\pi$ has trivial central character and $\dim\pi_{v}^{U_{v}}=1$ for $v$ finite. Explicitly, for $v\in S_{B}$, $\pi_{v}$ is the trivial representation of $B_{v}^{\times}$. Its Jacquet-Langlands correspondence to ${\mathrm{GL}}_{2}(F_{v})$ is the the Steinberg representation. In this cases, the root number $\epsilon(\pi_{v})=-1$. For $v\not\in S_{B}$, if $U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)$ for $v\not\in S_{B}$, then $\pi_{v}$ is an unramified representation of ${\mathrm{GL}}_{2}(F_{v})$. If $U_{v}=F_{v}^{\times}\langle{\Gamma_{0}\left({\mathfrak{m}}_{v}\right),g_{v}}\rangle$, then $\pi_{v}$ is the the Steinberg representation twisted by the unique nontrivial unramified quadratic character of $F_{v}^{\times}$. In both cases, the root number $\epsilon(\pi_{v})=1$.
@@
At an infinite place, $\epsilon(\pi_{v})=-1$. At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. And at every infinite place $v$, as we have seen above the lemma, $\epsilon(\pi_{v})=\epsilon(B_{v})$. So we have the first of the following equations $\epsilon(\pi)=-\prod_{v}\epsilon(B_{v})=-1.$
@@
If $v\in S_{B}$ so that $\pi_{v}=1$, the lemma is trivial. Otherwise, the lemma is [9, Proposition 6.1, 6.3]. ∎
@@
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$, then $\epsilon(\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
Let $\pi_{i,{\mathrm{f}}}$ denote the finite component of $\pi_{i}$. Then $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2). The main theorem of this section is the following.
@@
  • every geometrically connected component of $X_{U}$ is hyperelliptic.
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$, then
@@
The generalized Gross–Kudla conjecture [7] is proved in [24] under the first two assumptions in the theorem. By this and (5.4), the central derivative $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ is a multiple of the height of $\Delta_{3}]_{f_{1},f_{2},f_{3}}$, where Let $f_{i}\neq 0\in\pi_{i,{\mathrm{f}}}^{U}$, unique up to scalar. But by 4.5, $\Delta_{3}]_{f_{1},f_{2},f_{3}}=0$. Now the theorem follows from $5.$ ∎
@@ 6. Decomposable triple product $L$-functions
With the generalized Gross–Kudla conjecture[7, 24], we have used the (modified) diagonal cycle to study triple product $L$-functions. Now we take the opposite direction, and want to understand the behavior of the diagonal cycle for a general Shimura curve predicted by triple product $L$-functions. Keep the notations from the last section, except that now we abandon (5.1) and (5.2). We analyze two cases when the triple product $L$-function is decomposable.
First, if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual, then
@@
So if the root number $\epsilon(1/2,\pi)=-1$, then $L(s,\pi^{\otimes 3})$ has central vanishing order at least 2 at $s=1/2$. Thus by 5.3 and 5.5, 5.6 is trivial if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}.$
@@
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual and $\epsilon(1/2,\pi)=-1$, then $\Delta_{3}]_{f_{1},f_{2},f_{3}}=0$ for all $f_{i}\in\pi_{i,{\mathrm{f}}}^{U}$.
4.6 gives a special case of 6.1.
Second, if $\pi_{1}\simeq\pi_{2}^{\vee}=:\pi$, then
Assume $\pi_{3}$ is self-dual with root number $\epsilon(1/2,\pi_{3})=-1$. Then $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ has central vanishing order at least 3 at $s=1/2$ if and only if $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has central vanishing order at least 2 at $s=1/2$ or $L(s,\pi_{3})$ has central vanishing order at least 3 at $s=1/2.$ Then by 5.3, we have the following.
@@
In the situation of 5.6, if $\pi_{1}\simeq\pi_{2}=:\pi$ (which are self-dual), then $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has central vanishing order at least 2 at $s=1/2$ or $L(s,\pi_{3})$ has central vanishing order at least 3 at $s=1/2.$
Let $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2). For the diagonal cycle, if $\Delta_{3}]_{f_{1},f_{2},f_{3}}\neq 0$ in $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$, the generalized Gross–Kudla conjecture and the conjectural nondegeneracy of the height pairing imply that $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\neq 0$. Then the Beilinson–Bloch conjecture predicts that $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$ has dimension 1 and is thus generated by $\Delta_{3}]_{f_{1},f_{2},f_{3}}$. Moreover, by (6.1), $L^{\prime}(1/2,\pi_{3})\neq 0$ (assuming $\epsilon(1/2,\pi_{3})=-1$). Then by the general Gross–Zagier formula [11, 25] and 2.2 (which applies by 3.2 (2)), we propose the following conjecture.
Conjecture 6.3.
Assume $\pi_{1}\simeq\pi_{2}^{\vee}$, and $\pi_{3}$ is self-dual with $\epsilon(1/2,\pi_{3})=-1$. Let $(A,\iota)$ corresponding to $\pi_{3}$ under 3.1 (2). Then there exists a zero cycle $Z$ on $X_{U}$ such that $\Delta_{3}]_{f_{1},f_{2},f_{3}}$ is a multiple of $\Delta_{2}]_{f_{1},f_{2}}\times{\mathfrak{pr}}_{\iota,*}f_{3,*}[Z]$. Moreover, we can let $Z$ be a CM 0-cycle associated to some quadratic CM extension of $F$.
The relation between modified diagonal cycles and CM cycles was already explored in [2], as well as used [12, 15]. Our conjecture gives a complete explaination.

### v2

@@ 5. Triple product $L$-functions of central vanishing order at least 3
We keep the notations from the last section, except that from now on, the notation $\pi$ or $\pi_{i}$ will denote an automorphic representation of $B^{\times}$, instead of its finite part. We always assume that it is cuspidal holomorphic of weight $2$ in the sense that its Jacquet–Langlands correspondence to ${\mathrm{GL}}_{2,F}$ is a cuspidal automorphic representation holomorphic of weight $2$.
@@
Here the product is over all places of $F$. Let $\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})$ be the local root number of $\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v}$. A result of Prasad [17, 18] states the following.
@@
then it has dimension 1 and $\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=\epsilon(B_{v})$.
@@
for all finite places $v$ of $F$, then $\epsilon(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
@@
At every infinite place $v$, $\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-1$ (see [17, Section 9]). At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. So by 5.1, we have
$\epsilon(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=\prod_{v}\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-\prod_{v}\epsilon(B_{v})=-1.$
Let $S_{B}$ be the finite set of finite places $v$ of $F$ such that $B_{v}$ is a division quaternion algebra. Below we only consider a maximal open compact subgroup $U=\prod\limits_{v\text{ finite}}U_{v}\subset PB^{\times}({\mathbb{A}}_{\mathrm{f}})$ until the final remark (5.8). In other words, its $v$-component is
@@
where ${\mathfrak{m}}_{F_{v}}$ is the maximal ideal of ${\mathcal{O}}_{F_{v}}$, $\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}\right)$ is the subgroup of matrices in ${\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)$ that are upper triangular modulo ${\mathfrak{m}}_{F_{v}}$ and $g_{v}\in\begin{bmatrix}0&1\\ {\mathcal{O}}_{F_{v}}\backslash{\mathfrak{m}}_{F_{v}}&0\end{bmatrix}$. If $\pi^{U}\neq 0$, then $\pi$ has trivial central character and thus self-dual, and $\dim\pi_{v}^{U_{v}}=1$ for $v$ finite. Explicitly, for $v\in S_{B}$, $\pi_{v}$ is the trivial representation of $B_{v}^{\times}$. Its Jacquet-Langlands correspondence to ${\mathrm{GL}}_{2}(F_{v})$ is the the Steinberg representation. In this cases, the root number $\epsilon(1/2,\pi_{v})=-1$. For $v\not\in S_{B}$, if $U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)$ for $v\not\in S_{B}$, then $\pi_{v}$ is an unramified representation of ${\mathrm{GL}}_{2}(F_{v})$. If $U_{v}=F_{v}^{\times}\langle{\Gamma_{0}\left({\mathfrak{m}}_{v}\right),g_{v}}\rangle$, then $\pi_{v}$ is the the Steinberg representation twisted by the unique nontrivial unramified quadratic character of $F_{v}^{\times}$. In both cases, the root number $\epsilon(1/2,\pi_{v})=1$.
This lemma is not needed for the results in this section. We will use it later.
@@
At an infinite place, $\epsilon(1/2,\pi_{v})=-1$. At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. And at every infinite place $v$, as we have seen above the lemma, $\epsilon(1/2,\pi_{v})=\epsilon(B_{v})$. So we have the first of the following equations $\epsilon(1/2,\pi)=-\prod_{v}\epsilon(B_{v})=-1.$
This lemma is essential for the results in this section.
@@
If $v\in S_{B}$ so that $\pi_{v}=1$, the lemma is trivial. Otherwise, the lemma is [8, Proposition 6.1, 6.3]. ∎
@@
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$, then $\epsilon(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
Let $\pi_{i,{\mathrm{f}}}$ denote the finite component of $\pi_{i}$. Then $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2).
The main theorem of this section is the following.
@@
  • every geometrically connected component of $X_{U}$ is hyperelliptic, and
  • $\pi_{i}^{U}\neq 0$ for $i=1,2,3$.
If $\pi_{i}$’s are pairwise non-isomorphic, then
@@
The generalized Gross–Kudla conjecture [6] is proved in [24] under the first two assumptions in the theorem. By this and (5.3), the central derivative $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ is a multiple of the height of $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$, where $f_{i}\neq 0\in\pi_{i,{\mathrm{f}}}^{U}$, unique up to scalar. But by 4.5 (3), $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$. Now the theorem follows from $5.$ ∎
We will use this theorem to prove 1.1.
Replacing 4.5 (3) by 4.6 (2), we have the following more general theorem.
Theorem 5.7.
Assume
  • $U$ satisfies (5.1) and (5.2),
  • every geometrically connected component of $X_{U}$ is hyperelliptic, and
  • $\pi_{i}^{U}\neq 0$ for $i=1,2,3$,
  • $\pi_{3}\not\simeq\pi_{1}$, $\pi_{3}\not\simeq\pi_{2}$.
Let $f_{3}\neq 0\in\pi_{3,{\mathrm{f}}}^{U}$, unique up to scalar. If ${\mathfrak{pr}}_{f_{3},*}[{\mathfrak{e}}]=0$, then
${\mathrm{ord}}_{s=1/2}L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\geq 3.$
When $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}$, $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ has central vanishing order at least 3 automatically. We discuss this in the coming section.
Remark 5.8.
(1) The condition we impose on the levels of $\pi_{i}$’s, i.e., on $U$, above is for the purpose of getting 5.4, especially (5.3). Slightly more generally, we may further allow $U_{v}$ to be $F_{v}^{\times}{\mathcal{O}}_{B_{v}}^{\times}$ for $v\in S_{B}$ and the maximal order ${\mathcal{O}}_{v}$ of $B_{v}$, and allow $U_{v}$ to be $F_{v}^{\times}\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}\right)$ for $v\not\in S_{B}$. Then we need to match the representations so that the lemma still holds. For example, for $v\in S_{B}$, let $\pm 1$ be the two representations of $B_{v}^{\times}/F_{v}^{\times}O_{v}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$. Then we need the product of $\pi_{i,v}$’s to be 1. In particular, if the Atkin–Lehner involution, defined by the generator $B_{v}^{\times}/F_{v}^{\times}O_{v}^{\times}$, is hyperelliptic, then all $\pi_{i,v}$’s are 1-1. And the lemma will not hold.
(2) For $v\not\in S_{B}$, more generally, the Atkin–Lehner involution is defined if $U_{v}=F_{v}^{\times}\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}^{n_{v}}\right)$ for some positive integer $n_{v}$. Indeed, any $g_{v}\in\begin{bmatrix}0&1\\ {\mathfrak{m}}_{F_{v}}^{n-1}\backslash{\mathfrak{m}}_{F_{v}}^{n}&0\end{bmatrix}$ normalizes $U_{v}$ and defines the same Atkin–Lehner involution $w_{v}$ on $X_{U}$ at $v$. If $\dim\pi_{i,v}^{U_{v}}=1$, then $w_{v}$ acts on $\pi_{i,v}^{U_{v}}$ by $\pm 1$. If the action is by 1-1 and $f_{i}\neq 0\in\pi_{i,{\mathrm{f}}}^{U}$, one can use $w_{v}$ to show that $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$. Indeed, $(w_{v},w_{v},w_{v})$ fixes $\Delta_{3}]$ and thus fixes $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$, but sends each $f_{i}$ to $f_{i}$. The case when $w_{v}$ is a hyperelliptic involution is simply the case that any $\pi$ with $\pi_{i,{\mathrm{f}}}^{U}\neq 0$ satisfies that $\dim\pi_{i,v}^{U_{v}}=1$ and $w_{v}$ acts on $\pi_{i,v}^{U_{v}}$ by 1-1. However, in this case, (5.3) fails by the same argument. Thus $w_{v}$ being a hyperelliptic involution can not be used to produce higher central vanishing of $L$-functions.
(3) Finally, the above remarks on an Atkin–Lehner involution at a single finite place extend to a product of Atkin–Lehner involutions at several finite places in an obvious way.
@@ 6. Decomposable triple product $L$-functions
With the generalized Gross–Kudla conjecture[6, 24], we have used the (modified) diagonal cycle to study triple product $L$-functions. Now we take the opposite direction, and want to understand the behavior of the diagonal cycle for a general Shimura curve predicted by triple product $L$-functions. Keep the notations from the last section, except that now we abandon (5.1) and (5.2). We analyze two cases when the triple product $L$-function is decomposable. Such consideration appeared in an earlier version of [24] for the heights of the diagonal cycle.
First, if $\pi_{1}\simeq\pi_{2}^{\vee}=:\pi$, then
Assume $\pi_{3}$ is self-dual with root number $\epsilon(1/2,\pi_{3})=-1$. Then $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ has vanishing order at least 3 at $s=1/2$ if and only if $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has vanishing order at least 2 at $s=1/2$ or $L(s,\pi_{3})$ has vanishing order at least 3 at $s=1/2.$ Then by 5.3, we have the following.
Proposition 6.1.
In the situation of 5.7, if $\pi_{1}\simeq\pi_{2}=:\pi$ (which are self-dual) and $L^{\prime}(1/2,\pi_{3})\neq 0$, then $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has vanishing order at least 2 at $s=1/2$.
Now, we consider the diagonal cycle. Let $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2). If $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}\neq 0$ in $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$, the generalized Gross–Kudla conjecture and the conjectural nondegeneracy of the height pairing imply that $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\neq 0$. Then the Beilinson–Bloch conjecture predicts that $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$ has dimension 1 and is thus generated by $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$. Moreover, by (6.1), $L^{\prime}(1/2,\pi_{3})\neq 0$ (assuming $\epsilon(1/2,\pi_{3})=-1$). Then by the general Gross–Zagier formula [10, 25] and 2.2 (which applies by 3.2 (2)), we propose the following conjecture.
Conjecture 6.2.
Assume $\pi_{1}\simeq\pi_{2}^{\vee}$, and $\pi_{3}$ is self-dual with $\epsilon(1/2,\pi_{3})=-1$. Let $(A,\iota)$ corresponding to $\pi_{3}$ under 3.1 (2). Then there exists a zero cycle $Z$ on $X_{U}$ such that $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$ is a multiple of $\Delta_{2}]_{f_{1}\otimes f_{2}}\times{\mathfrak{pr}}_{\iota,*}f_{3,*}[Z]$. Moreover, we can let $Z$ be a CM 0-cycle associated to some quadratic CM extension of $F$.
The relation between modified diagonal cycles and CM cycles was already explored in [1], as well as used [12, 15]. Our conjecture gives a complete explanation. Alternatively, in an earlier version of [24], the authors proposed to use an ample and symmetric line bundle associated to the polarization $A_{1}\to A_{2}\simeq A_{1}^{\vee}$ to construct 0-cycles that replace CM cycles in this case. See also [28, 5.3].
Second, if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual, then
@@
So if the root number $\epsilon(1/2,\pi)=-1$, then $L(s,\pi^{\otimes 3})$ has vanishing order at least 2 at $s=1/2$. Thus by 5.3 and 5.5, 5.6 is trivial if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}.$
@@
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual and $\epsilon(1/2,\pi)=-1$, then $\Delta_{3}]_{f}=0$ for all $f\in\boxtimes_{i=1}^{3}\pi_{i}^{U}$.
2.1 (which applies by 3.2 (2)) gives the following proposition. It contains a special case of 6.3.
@@
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$, for $f\in{\mathrm{sym}}^{3}\pi$, $\Delta_{3}]_{f}=0$.
Remark 6.5.
If $\epsilon(1/2,\pi)=1$, the vanishing in the proposition follows from 4.2 and a result of Prasad [17, Theorem 6]. This is purely local and representation-theoretical.
7. Hyperelliptic modular curves
Consider the case of modular curves, i.e., let $B={\mathrm{M}}_{2}({\mathbb{Q}}).$ A Shimura curve $X_{U}$ with $U$ satisfying (5.2) is the quotient $X_{0}^{*}(N)$ of the classical modular curve $X_{0}(N)$ by the full Atkin–Lehner group, for some $N$ square free. (See 5.8 (2) for an adelic definition of Atkin–Lehner involutions.) Let $U(N)\subset{\mathrm{PGL}}_{2}({\mathbb{A}}_{\mathrm{f}})$ be the corresponding open compact subgroup.
By [11], the list of square free $N$’s with $X_{0}^{*}(N)$ hyperelliptic consists of the following 39 numbers:
[…]
Moreover, these $X_{0}^{*}(N)$’s all have genus 2. Thus we can not apply 5.6. Instead, we want to apply 5.7. We need the following lemma.
Lemma 7.1.
Let $N$ be square free, Assume
  • (1)
    there exists a prime factor $p\mid N$ that is inert in $\mathbb{Q}(i)$, and
  • (2)
    there exists a (possibly equal) prime factor $q\mid N$ that is inert in $\mathbb{Q}(\sqrt{-3})$.
For a cuspidal automorphic representation $\pi$ of ${\mathrm{GL}}_{2,{\mathbb{Q}}}$ holomorphic of weight $2$, if $f\neq 0\in\pi^{U(N)}$, then ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0$ on $X_{0}^{*}(N)$.
Proof.
The divisor of elliptic points has a decomposition ${\mathfrak{e}}={\mathfrak{e}}_{2}+{\mathfrak{e}}_{3}$ into a CM divisor with CM by $\mathbb{Q}(i)$ and $\mathbb{Q}(\sqrt{-3})$. We show ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}_{2}]={\mathfrak{pr}}_{f,*}[{\mathfrak{e}}_{3}]=0$. We prove the first equation and the proof of the second is the same.
Let $E=\mathbb{Q}(i)$ embedded in $B={\mathrm{M}}_{2}({\mathbb{Q}})$, uniquely up to conjugation. Using the formulation in [25, 1.3.2], we only need to show ${\mathrm{Hom}}_{E_{p}^{\times}}(\pi_{p},{\mathbb{C}})=0.$ Indeed, $\pi_{p}$ is the the Steinberg representation twisted by the unique nontrivial unramified quadratic character of ${\mathbb{Q}}_{p}^{\times}$. Let $D_{p}$ be the unique division quaternion algebra over ${\mathbb{Q}}_{p}$ and $E_{p}$ embedded in $D_{p}$, uniquely up to conjugation. The Jacquet–Langlands correspondence of $\pi_{p}$ to $D_{p}^{\times}$ is the unique nontrivial (quadratic) character $\chi$ of $D_{p}^{\times}/{\mathbb{Q}}_{p}^{\times}O_{p}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$, where $O_{p}$ is a maximal order of $D_{p}$. Since ${\mathrm{Hom}}_{E_{p}^{\times}}(\chi,{\mathbb{C}})\neq 0$ by assumption (1), by the Tunnell–Saito dicohotmy [25, Theorem 1.3], ${\mathrm{Hom}}_{E_{p}^{\times}}(\pi_{p},{\mathbb{C}})=0.$ The lemma is proved. ∎
In $\mathbb{Q}(i)$, a prime $p$ is inert if and only if $p\equiv 3\pmod{4}$. In $\mathbb{Q}(\sqrt{-3})$, a prime $q$ is inert if and only if $q\equiv 2\pmod{3}$. Among the $39$ squarefree $N$ in the list, the following $26$ contain at least one prime factor $p\equiv 3\pmod{4}$ and at least one prime factor $q\equiv 2\pmod{3}$.
Table 1. Squarefree NN in Equation 7.1 with prime factors inert in $\mathbb{Q}(i)$ and prime factors inert in $\mathbb{Q}(\sqrt{-3})$.
[…]
Theorem 7.2.
For $N$ in Table 1, let $f,g$ be (the only) two distinct newforms of level $\Gamma_{0}(N)$ and Atkin–Lehner sign +1+1 at all primes dividing $N$. Then
${\mathrm{ord}}_{s=2}L\left(s,{\mathrm{sym}}^{2}f\otimes g\right)\geq 2.$
Proof.
By 7.1, 5.7 applies. By 6.1, we only need to show that $L^{\prime}(1,g)\neq 0$. This can be read from LMFDB [14]. ∎